OpenStudy (anonymous):

find the oblique asymptotes of f(x) = (9x-x^3)/(x^2-4). Is this oblique asymptote a linear or nonlinear asymptote

OpenStudy (amistre64):

the bottom and top exponent differ by only one degree which means that this oblique is linear....

OpenStudy (amistre64):

better put... that highest exponents on top and bottom only differ by one degree; so the oblique is a line

OpenStudy (anonymous):

Can you explain how I could solve the equation

OpenStudy (amistre64):

solve it for the oblique you mean?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

long division: -x ------------ x^2 -4 |-x^3 +9x x^3 -4 --------- 9x -4 <- remainder the oblique takes the shape of: 9x-4 -x + ------ x^2-4 the left side part goes to zero is x gets large which means that y = -x+0 or simply y = -x is the equation of the line for your oblique asymptote

OpenStudy (amistre64):

and by left side I mean right side.... Oy!!

OpenStudy (anonymous):

Ok, let me study this for a second. thanks

OpenStudy (amistre64):

your seconds up :)

OpenStudy (amistre64):

as the bottom of a fraction gets very large: 1 ----------------------------- 1000000000000000000000000 the value of the fraction get very very very tiny and gets close to zero at the ends.

OpenStudy (amistre64):

so anything with an "x" in the bottom gets thrown out and counted for zero

OpenStudy (anonymous):

y = -x+0 would mark the end of the equation?